Suppose that a cial bank wants to buy Treasury bills These instruments pay $4.000 in one yoar and are currently selling The yield to maturity of these bonds is % (Round your response to two dec mal places)
Answer #1

Yield = (Par value - price) / Price*100

where: Par value = $4000

Current value = $4250

Yield = (4000-4250/4250)*100 = -(250/4250*100) = -5.88%

Hence the bank suffers a loss in the purchase of the T bill